# New GPU developments in the Madgraph CUDACPP plugin: kernel splitting, helicity streams, cuBLAS color sums

### v1

@@ 3.1 Helicity streams (“ihel1”)
The internal substructure of the sigmaKin ME engine in the current version (ihel0) of CUDACPP is illustrated schematically in the top-left diagram in Fig. 2. This gives a visual representation of the pseudo-code in Table 1. As mentioned previously and as visible in the diagram, the two main components of the calculation for each event, namely the computation of wavefunctions and dual amplitudes for a given color flow from Feynman diagrams, and their squaring and sum over all color flows, are performed sequentially for all helicities. This workflow makes it impossible to split these two components into separate kernels, because the loop over helicities is effectively inside the loop over events.
@@
The very first step in my kernel splitting developments was therefore, quite naturally, to reverse this situation and make the event loop the innermost loop, inside an outermost loop on helicities. This was the focus of my work with Olivier Mattelaer during the 2022 GPU hackathon (and is quite likely an idea that he suggested, in fact). In practice, the main change to achieve here was to turn sigmaKin from a __global__ device kernel into a host function, and to turn instead calculate_wavefunction from a __device__ function callable by a kernel into a kernel itself. Other computations also had to be modified: for instance, the selections of event-by-event colors and helicities, which the sigmaKin kernel was performing outside the helicity loop (see Table 1) have now also been turned into separate GPU kernels. These changes, which I achieved in November 2024 bib:ks2024 , by themselves already provide a moderate increase of throughputs for small grids.
@@
The real breakthrough, however, came when the parallel calculations for different helicities were moved to separate CUDA Streams. In this “ihel1” version of the software, which is illustrated in the top-right diagram in Fig. 2, ME throughputs on NVidia GPUs reach their peak performance with much smaller grids than in the current ihel0 version, for all physics processes and floating precisions I tested. The peak throughput themselves are also increased by around 10-20%. This can be seen in Fig. 4, by comparing the blue (ihel0) and orange (ihel1) curves. The improvement is especially impressive for complex processes like $gg\!\rightarrow\!t\bar{t}ggg$, where peak throughputs are reached with $\cal O$}(100) events per grid in ihel1, as opposed to $\cal O$}(10k) in the current ihel0. These improvements from ihel1 are essentially the single most important progress described in this paper. Most likely, the improvement comes from the increase in parallelism of the workflow: instead of a single kernel launch in each offloading cycle, which runs for a long time because it internally loops over helicities, there are now several, much shorter kernels launched in parallel, one for each helicity, in a separate GPU stream for each helicity.
@@
For AMD GPUs, where the same solution was implemented using HIP Streams, the benefits are much less clear, as shown in Fig. 5: throughputs increase with both small and large grids for complex processes like $gg\!\rightarrow\!t\bar{t}gg$, but for simpler processes the opposite effect is observed. The code refactoring in ihel1 was also propagated to the SIMD C++ code, and tested on a reference Intel node, which has excellent AVX512 performance thanks to the presence of 2 FMA units. I will not give detailed numbers, but my tests show that the throughputs are essentially unchanged for all SIMD builds, within the expected 1-2% fluctuations.
@@ 3.2 Color sum as a separate GPU kernel (“ihel2”)
The next step in my kernel splitting developments consisted, quite naturally, in separating the two main components of the ME calculation, namely, (1) the calculation from Feynman diagrams of the dual amplitudes for ncolor color flows (jamp: a vector $J$ with ncolor complex elements) and (2) the color sum. Initially, I simply split the calculate_wavefunction kernel, which was doing both computations in ihel1, into two separate kernels: calculate_jamps, which calculates the jamp $J$, and color_sum, which computes the quadratic form $J^{H}(C)J$ for a symmetric real color matrix $C$ using the vector $J$ and its conjugate transpose $J^{H}$. This “ihel2” version of the software is illustrated in the bottom-left diagram in Fig. 2. One important difference in this case is that the jamp variable is no longer a local variable inside calculate_wavefunction, as it was in ihel0 (see Table 1) and ihel1: instead, it is now a GPU global memory buffer that is allocated outside the MEK component and is accessed, with the appropriate layout decoding provided by a new class DeviceAccessJamps, in both the calculate_jamps and color_sum kernels.
As can be seen in Fig. 4, by comparing the orange (ihel1) and green (ihel2) curves, on Nvidia GPUs this further change in the software yields an additional increase of the peak throughput by 10-20% for complex processes like $gg\!\rightarrow\!t\bar{t}gg$ and $gg\!\rightarrow\!t\bar{t}ggg$. For simpler processes, conversely, it results in a minor decrease of peak throughputs; in my opinion, this is a moderate cost that can be tolerated, as speeding up complex processes is more important. On AMD GPUs, as can be seen in Fig. 5, the additional change results in a minor increase in peak throughputs for $gg\!\rightarrow\!t\bar{t}gg$, but for simpler processes ihel2 is almost indistinguishable from ihel1. In the CPU implementation, throughputs are again essentially unchanged for all SIMD builds within 1-2%.
@@ 3.3 Color sum dispatcher to kernel or BLAS (“ihel3”)
    color_sum_kernel. By default, the calculation is performed using a kernel color_sum_kernel. This is essentially the same code as in the ihel2 software, with one minor difference: the GPU global memory layout of the dual amplitudes, which was an SOA jamp[ncolor][2][nevt] in ihel2, is now an SOA jamp[2][ncolor][nevt], because this makes it easier to separate the real and imaginary parts of the dual amplitudes for the BLAS calculation, and the same layout is used for simplicity in both the kernel and BLAS implementations. As can be seen in Fig. 4 by comparing the green (ihel2) and red (ihel3) curves, the achieved performance is indistinguishable from that of ihel2.
@@
The performances of the cuBLAS and kernel implementation of color sums on NVidia GPUs are compared in Fig. 4, where they are referred to as “ihel3b” (purple) and “ihel3” (red), even if they both result from the same ihel3 code, with and without an environment variable set at runtime. The picture clearly shows that BLAS performs much worse than CUDA kernels for the simpler $g\!\rightarrow\!t\bar{t}, gg$$g\!\rightarrow\!t\bar{t}g$ and $gg\!\rightarrow\!t\bar{t}gg$ physics processes. For the more complex $gg\!\rightarrow\!t\bar{t}ggg$ process, the situation is more complex: for small grids, the kernel implementation is faster for all floating point precisions; for large grids, however, the BLAS implementation eventually becomes as fast as the kernel implementation, and in double and float precision (but not in mixed precision) it is eventually faster for very large grids. This is interesting, but of not much practical relevance, as most production workflows would use small grids to keep event generation jobs more manageable. On AMD GPUs, Fig. 5 shows that the kernel version of color sums is always much better than the corresponding hipBLAS implementation. Keeping all these observations into account, it seems appropriate to compute color sums using kernels instead of BLAS by default.
Figure 6: Time spent in the color sum as a fraction of the total time spent in the sigmaKin ME engine, for different physics processes and software configurations, in the ihel3 version of the software. The 3 plots correspond to 3 floating point precisions. Measurements performed on a node equipped with an NVidia V100 GPU and an Intel Xeon Silver 4216 CPU (which supports AVX512 but gives better results for CUDACPP in the “512y” mode than in the “512z” mode because it has only one FMA unit). The results shown for CUDA are based on the default ihel3 implementation of color sums using GPU kernels.
Figure 7: Time spent in the color sum as a fraction of the total time spent in the sigmaKin ME engine for the $gg\!\rightarrow\!t\bar{t}ggg$ process, in the CUDA kernel (ihel3) and CUDA/BLAS (ihel3b) implementations, as a function of the number of events nevt processed in parallel. The 3 plots correspond to 3 floating point precisions. For the CUDA kernel implementation, nevt is the product of the number of threads per block, which is fixed to 32, by the varying number of blocks in the GPU grid. For the CUDA/BLAS implementation, nevt is the number of events in the dual amplitude vectors on which the BLAS color matrix multiplication is computed. Measurements performed for an NVidia V100 GPU, on the same node used to prepare Fig. 6.
@@ 3.3.1 Color sum profiling
The results of my analysis, which are shown in Fig. 6 for the CUDA and SIMD C++ backends, are somewhat surprising: for $gg\!\rightarrow\!t\bar{t}ggg$, in the ihel3 version of the software, the color sum implementation in CUDACPP represents only 5 to 10% of the total time to compute the ME in all SIMD CPU modes, while in the CUDA kernel implementation this fraction is around 28%, 15% and 6% in double, single and mixed precision modes, respectively. These results are particularly good for the mixed precision mode, as they indicate that our previous optimizations have managed to reduce the time taken by the color sum to a level where this is no longer a bottleneck of the ME calculation. In retrospective, as a consequence, the benefits that one may hope to obtain on GPUs from BLAS are quite limited.
In Fig. 6, the results for BLAS color sums are not shown: this is because they heavily depend on the number of events nevt processed in each GPU cycle, which corresponds to the size of the event vector in BLAS and to the GPU grid size for the calculate_jamp kernels. This is illustrated in Fig. 7, which shows that the efficiency of BLAS color sums increases as nevt increases, while that of the kernel implementation is essentially constant. This plot is consistent with the trend shown in Fig. 4 and is useful to better understand it.
In the C++ implementation, one reason for the relatively low time footprint of the color sum may be a set of optimizations that I introduced in October 2022 and never documented in detail. To start with, as discussed above for BLAS in Eq. 1, I simplified the color sum on the complex vector of dual amplitudes as the sum of two quadratic forms on real vectors. In addition, I also rewrote the multiplication as one involving a triangular matrix, to ensure that the non-diagonal terms of the symmetric matrix are only used once in the calculation.
@@ 3.3.2 Tensor Core profiling
As the use of Tensor Cores was one of the main motivations for testing the use of cuBLAS for color sums, in Table 3 I present the results of some quick studies using the NVidia NSight Compute (ncu) profiler. Unlike the other results shown in this paper, these were obtained using a more recent NVidia A100 GPU, which has more advanced Tensor Core features than a V100. The results of this test, which was performed using CUDACPP ihel3 with mixed floating precision, clearly show that Tensor Cores are used by cuBLAS color sums. In particular, the cublasSgemm multiplication $(C)V$ makes significant use of Tensor Cores, but only if TF32 math mode is explicitly enabled, as otherwise cuBLAS seems to prefer traditional SIMT kernels. As for the multiplication of $V^{t}$ by $(C)V$ in cublasSgemmStridedBatched, two different implementations are used by cuBLAS for small and large numbers of events, but neither of them makes a significant use of Tensor Cores. Other tests in different configurations, whose results are not presented here in detail, indicate that cuBLAS color sums make a significant use of Tensor Cores (irrespective of whether TF32 math mode is enabled or not) when double precision is used. This is somewhat surprising, as Tensor Cores are designed and optimized for lower precision calculations in AI. More detailed studies could be useful to understand the usefulness of cuBLAS for CUDACPP color sums on more recent NVidia GPUs, with much more advanced Tensor Core features.
3.4 Feynman diagrams as separate kernels (“ihel4”)

### v2

@@ 3.1 Helicity streams (“ihel1”)
The internal substructure of the sigmaKin ME engine in the previous version of CUDACPP before my kernel splitting changes (ihel0) is illustrated schematically in the top-left diagram in Fig. 3. This gives a visual representation of the pseudo-code in Table 1. As mentioned previously and as visible in the diagram, the two main components of the calculation for each event, namely the computation of wavefunctions and dual amplitudes for a given color flow from Feynman diagrams, and their squaring and sum over all color flows, are performed sequentially for all helicities. This workflow makes it impossible to split these two components into separate kernels, because the loop over helicities is effectively inside the loop over events.
@@
The very first step in my kernel splitting developments was therefore, quite naturally, to reverse this situation and make the event loop the innermost loop, inside an outermost loop on helicities. This was the focus of my work with Olivier Mattelaer during the 2022 GPU hackathon (and is quite likely an idea that he suggested, in fact). In practice, the main change to achieve here was to turn sigmaKin from a __global__ device kernel into a host function, and to turn instead calculate_wavefunction from a __device__ function callable by a kernel into a kernel itself. Other computations also had to be modified: for instance, the selections of event-by-event colors and helicities, which the sigmaKin kernel was performing outside the helicity loop (see Table 1) have now also been turned into separate GPU kernels. As discussed in Ref. bib:ks2024 , by themselves these changes already provide a moderate increase of throughputs for small grids.
@@
The real breakthrough, however, came when the parallel calculations for different helicities were moved to separate CUDA Streams. In this “ihel1” version of the software, which is illustrated in the top-right diagram in Fig. 3, ME throughputs on NVidia GPUs reach their peak performance with much smaller grids than in the previous ihel0 version, for all physics processes and floating precisions I tested. The peak throughput themselves are also increased by around 10-20%. This can be seen in Fig. 4, by comparing the blue (ihel0) and orange (ihel1) curves. The improvement is especially impressive for complex processes like $gg\!\rightarrow\!t\bar{t}ggg$, where peak throughputs are reached with $\cal O$}(100) events per grid in ihel1, as opposed to $\cal O$}(10k) in the current ihel0. These improvements from ihel1 are essentially the single most important progress described in this paper. Most likely, the improvement comes from the increase in parallelism of the workflow: instead of a single kernel launch in each offloading cycle, which runs for a long time because it internally loops over helicities, there are now several, much shorter kernels launched in parallel, one for each helicity, in a separate GPU stream for each helicity.
@@
For AMD GPUs, where the same solution was implemented using HIP Streams, the benefits are much less clear, as shown in Fig. 5: throughputs increase with both small and large grids for complex processes like $gg\!\rightarrow\!t\bar{t}gg$, but for simpler processes the opposite effect is observed [^5] I show no results for $gg\!\rightarrow\!t\bar{t}ggg$ on AMD GPUs, because the code failed to build on the system I used (“error: unhandled SGPR spill to memory”) with the ihel0 software. It is possible that this issue could be addressed using the Feynman diagram splitting techniques in ihel5-ihel6, but I did not attempt this. More generally, I stress that, while in this paper I present results also for AMD GPUs, I never attempted any serious optimization targeting them specifically, and it is perfectly possible that CUDACPP performance on this platform could be significantly improved with relatively little effort. . The code refactoring in ihel1 was also propagated to the C++ code, and tested on a reference Intel CPU; these showed that the throughputs are essentially unchanged for all SIMD builds.
@@ 3.2 Color sum as a separate GPU kernel (“ihel2”)
The next logical step in my kernel splitting developments consisted in separating the two main components of the ME calculation, namely, (1) the calculation from Feynman diagrams of the dual amplitudes for ncolor color flows (jamp: a vector $J$ with ncolor complex elements) and (2) the color sum. Initially, I simply split the calculate_wavefunction kernel, which was doing both computations in ihel1, into two separate kernels: calculate_jamps, which calculates the jamp $J$, and color_sum, which computes the quadratic form $J^{H}(C)J$ for a symmetric real color matrix $C$ using the vector $J$ and its conjugate transpose $J^{H}$. This “ihel2” version of the software is illustrated in the bottom-left diagram in Fig. 3. One important difference in this case is that the jamp variable is no longer a local variable inside calculate_wavefunction, as it was in ihel0 (see Table 1) and ihel1: instead, it is now a GPU global memory buffer that is allocated outside the MEK component and is accessed, with the appropriate layout decoding provided by a new class DeviceAccessJamps, in both the calculate_jamps and color_sum kernels.
As the plots I prepared are already packed, I decided not to present the ihel2 results explicitly: for reasons that will become clear in the following, in fact, these are essentially identical to the “ihel3” results shown by the solid-red curve in Fig. 4. As can be seen by comparing this to the orange curve for ihel1, on Nvidia GPUs the further change in the software from the split of Feynman diagrams and color sums yields an additional increase of the peak throughput by 10-20% for complex processes like $gg\!\rightarrow\!t\bar{t}gg$ and $gg\!\rightarrow\!t\bar{t}ggg$. For simpler processes, conversely, it results in a minor decrease of peak throughputs; in my opinion, this is a moderate cost that can be tolerated, as speeding up complex processes is more important. On AMD GPUs, as shown in Fig. 5, the additional change results in a minor increase in peak throughputs for $gg\!\rightarrow\!t\bar{t}gg$, but for simpler processes ihel2 is almost indistinguishable from ihel1. In the CPU implementation, throughputs are again essentially unchanged for all SIMD builds.
@@ 3.3 Color sum dispatcher to kernel or BLAS (“ihel3”)
Figure 7: Time spent in the color sum as a fraction of the total time spent in the sigmaKin ME engine, for different physics processes and software configurations, in the ihel3 version of the software. Lower is better. The 3 plots correspond to 3 floating point precisions. Measurements performed on a node equipped with an NVidia V100 GPU and an Intel Xeon Silver 4216 CPU (which supports AVX512 but gives better results for CUDACPP in the “512y” mode than in the “512z” mode because it has only one FMA unit). The results shown for CUDA are based on the default ihel3 implementation of color sums using GPU kernels.
@@
    color_sum_kernel. By default, the calculation is performed using a kernel color_sum_kernel. This is essentially the same code as in the ihel2 software, with one minor difference: the GPU global memory layout of the dual amplitudes, which was an SOA jamp[ncolor][2][nevt] in ihel2, is now an SOA jamp[2][ncolor][nevt], because this makes it easier to separate the real and imaginary parts of the dual amplitudes for the BLAS calculation, and the same layout is used for simplicity in both the kernel and BLAS implementations. The achieved performance is represented by the “ihel3” solid-red curve in Fig. 4. As mentioned above, these results are indistinguishable from those of ihel2, which I therefore decided not to show this explicitly to simplify the plot; the comparison, however, is available in my previous preprint of this article bib:preprintV1 .
@@
The performances of the cuBLAS and kernel implementation of color sums on NVidia GPUs using the ihel3 codebase are compared in Fig. 4, where they are referred to as “ihel3b” (dashed-red) and “ihel3” (solid-red), the only difference between them being an environment variable set at runtime. The picture clearly shows that the BLAS implementation in the ihel3 codebase performs much worse than CUDA kernels for the simpler $g\!\rightarrow\!t\bar{t}, gg$$g\!\rightarrow\!t\bar{t}g$ and $gg\!\rightarrow\!t\bar{t}gg$ physics processes. For the more complex $gg\!\rightarrow\!t\bar{t}ggg$ process, the situation is less clear-cut: for small grids, the kernel implementation is faster for all floating point precisions; for large grids, however, the ihel3 BLAS implementation eventually becomes as fast as the kernel implementation, and in double and float precision (but not in mixed precision) it is eventually faster for very large grids. This is interesting, but of not much practical relevance, as most production workflows would use small grids to keep event generation jobs more manageable. This observation, however, was the starting point of my further refinement of BLAS color sums in the “ihel3p1” version of the code, described in a later subsection. On AMD GPUs, Fig. 5 shows that the kernel version of color sums is always much better than the corresponding hipBLAS implementation, in the ihel3 codebase.
Figure 8: Time spent in the color sum as a fraction of the total time spent in the sigmaKin ME engine for the $gg\!\rightarrow\!t\bar{t}ggg$ process, as a function of the number of events nevt processed in parallel in one GPU iteration. Lower is better. In each plot, four curves are shown, using the ihel3 (red) and ihel3p1 (black) codebases, in the CUDA kernel (solid) and CUDA/BLAS (dashed) implementations. The 3 plots correspond to 3 floating point precisions. The handling of the nGoodHel good helicities is different in the various scenarios. In the two CUDA kernel implementations, nGoodHel kernels are launched in separate helicity streams, each one on a GPU grid of size nevt: in particular, nevt is the product of the number of threads per block, which is fixed to 32, by the varying number of blocks in the GPU grid. For the CUDA/BLAS implementations: in ihel3b, nGoodHel separate BLAS color matrix multiplications are computed, in different helicity streams, using dual amplitude vectors with nevt events; in ihel3b, a single BLAS color matrix multiplications is computed, in the default stream, using dual amplitude vectors with nGoodHel*nevt events. Measurements performed for an NVidia V100 GPU, on the same node used to prepare Fig. 7.
@@ 3.3.1 Color sum profiling
The results of my analysis, which are shown in Fig. 7 for the CUDA and SIMD C++ backends, are somewhat surprising: for $gg\!\rightarrow\!t\bar{t}ggg$, in the ihel3 version of the software, the color sum implementation in CUDACPP represents only 5 to 10% of the total time to compute the ME in all SIMD CPU modes, while in the CUDA kernel implementation this fraction is around 28%, 6% and 15% in double, mixed and single precision modes, respectively. These results are particularly good for the mixed precision mode, and indicate that our previous optimizations have managed to reduce the time taken by the color sum to a level where this is no longer a bottleneck of the ME calculation.
In Fig. 7, the results for BLAS color sums are not shown: this is because they heavily depend on the number of events nevt processed in each GPU cycle, which corresponds to the size of the event vector in BLAS and to the GPU grid size for the calculate_jamp kernels. These results are instead shown by the two red curves in Fig. 8, which show that the efficiency of BLAS color sums increases as nevt increases, while that of the kernel implementation is essentially constant. This plot is consistent with the trend shown in Fig. 4 and is useful to better understand it. The two other black curves in Fig. 8 refer to the further improvements in the “ihel3p1” version of the software, described further below.
In the C++ implementation, one reason for the relatively low time footprint of the color sum may be a set of optimizations that I introduced in October 2022 after some discussions in the team (and notably some suggestions by Olivier Mattelaer), and never documented in detail. To start with, as discussed above for BLAS in Eq. 1, I simplified the color sum on the complex vector of dual amplitudes as the sum of two quadratic forms on real vectors. In addition, I also rewrote the multiplication as one involving a triangular matrix, to ensure that the non-diagonal terms of the symmetric matrix are only used once in the calculation, with a factor 2. While writing this paper, I realized that this optimization had been applied to C++ code, but not yet to color sum kernels on GPUs: this is another point that I later fixed in the “ihel3p1” version, described further below.
@@ 3.3.2 Tensor core profiling
As the use of tensor cores was one of the main motivations for prototyping cuBLAS for color sums, I present in Table 3 the results of some quick studies using the NVidia NSight Compute (ncu) profiler. Unlike the results shown so far in this paper, these were obtained using an NVidia A100 GPU, which has more advanced tensor core features than the older V100 model. The results of this test, which was performed using CUDACPP ihel3 with mixed floating precision, clearly show that tensor cores are used by cuBLAS color sums. In particular, the cublasSgemm multiplication $(C)V$ makes significant use of tensor cores, but only if TF32 math mode is explicitly enabled, as otherwise cuBLAS seems to prefer traditional SIMT kernels. As for the multiplication of $V^{t}$ by $(C)V$ in cublasSgemmStridedBatched, two different implementations are used by cuBLAS for small and large numbers of events, but neither of them makes a significant use of tensor cores. Other tests in different configurations, whose results are not presented here in detail, indicate that cuBLAS color sums make a significant use of tensor cores (irrespective of whether TF32 math mode is enabled or not) when double precision is used. This is somewhat surprising, as tensor cores are designed and optimized for lower precision calculations in AI. A more detailed investigation of tensor cores in the CUDACPP cuBLAS implementation is beyond the scope of this paper, but could be useful for instance in the context of the HEP-SCORE benchmarking project. In particular, further investigations could be useful to understand the usefulness of cuBLAS for CUDACPP color sums on more recent NVidia GPUs, with much more advanced tensor core features.
3.4 Further improvements of color sums (“ihel3p1”)
As mentioned above, while writing the first preprint for this paper bib:preprintV1 I realised that two optimizations of color sums on GPUs could still be attempted. I have implemented both of them in a later version “ihel3p1” of the code, the results for which are represented by the black-solid (ihel3p1, BLAS disabled) and black-dashed lines (ihel3p1b, BLAS enabled) in Figures 4 and 5 for NVidia and AMD GPUs. These patches have already been merged in production with the rest of my first batch of changes to the code: the black lines in the figures, in other words, represent the performance of the current latest release CUDACPP v1.01.01. As shown in Fig. 6, I also note that the C++ implementation in this ihel3p1 release has essentially the same performance as the previous ihel0 release for all SIMD builds.
The first and simplest optimization was that, a previously done in C++, I rewrote the color sum kernel using a triangular matrix, to ensure that the non-diagonal terms of the symmetric matrix are only used once in the calculation with a factor 2, rather than twice. As shown by the bottom-left and bottom-right plots in Fig. 4, this small change does provide a minor but visible increase in peak throughput in $gg\!\rightarrow\!t\bar{t}ggg$, for double and single precision, but not for mixed precision where the time footprint of the color sum was already small. This is also confirmed by comparing the solid-red (ihel3) and solid-black (ihel3p1) curves in Fig. 8, which shows that the fraction of time spent in color sum kernels is reduced by around 5% in both double and single precision.
The second optimization, conversely, concerns BLAS and was motivated by the observations, discussed above and based on Figures 4 and 8, that BLAS throughputs significantly increase when large vectors are used. The basic idea, therefore, was to group together the color sums for all helicities as a single BLAS multiplication, rather than perform a separate BLAS multiplication for each helicity. The changes that I implemented are therfore the following: first, I removed the color sum dispatcher from the helicity loop, implementing a single dispatcher for all helicities rather than one per helicity; in the default case where color sum kernels are used, the dispatcher takes care of launching separate kernels for the different helicities, in the corresponding helicity streams; if BLAS is selected at runtime, however, the dual amplitudes from the different helicities are concatenated to form a single much larger array of events, as if the different helicities for each event were treated as different events, and a single BLAS multiplication is performed in the default stream. This is represented in the lower diagrams of Fig. 11 (this refers to a later ihel6 version of the software, which however only differs from ihel3p1 in the handling of Feynman diagrams). As shown in Fig. 4, the BLAS throughputs of this improved ihel3p1b implementation (dashed-black) are much higher than those of the previous ihel3b implementation (dashed-red). For $gg\!\rightarrow\!t\bar{t}ggg$ (bottom row), in particular, the performance receives a massive boost: BLAS color sums are now better than kernels in double precision, and have equivalent performance in mixed and single precision. This can also be seen in Fig. 8, which shows that the fraction of time spent in the color sum in $gg\!\rightarrow\!t\bar{t}ggg$ is now consistently in the 5–10% range when BLAS is used, independently of the number of events in each GPU iteration. For simpler processes, however, color sum kernels still maintain a higher performance: keeping this into account, it seems appropriate to keep color sum kernels as the default and BLAS as an optional alternative.
Figure 9: Schematic representation of the CUDACPP ME engine (sigmaKin), in the fourth kernel splitting scenario described in this paper: (ihel4) Feynman diagrams as individual kernels.
4 Kernel splitting II: Feynman diagram groups
After implementing helicity-level parallelism, separating Feynman diagrams from color sums and porting the latter to BLAS, the next step in my kernel splitting developments consisted, quite naturally, in trying to further decompose the calculate_jamps kernel, which in ihel3 computes dual amplitudes from all Feynman diagrams, into many shorter kernels. My initial motivation for attempting this was to investigate whether this may lead to further throughput increases, possibly thanks to the lower number of registers used by smaller GPU kernels (lower “register pressure”): I soon realised, however, that I was not able to achieve further throughput increases, because moderately small Feynman diagram kernels are still quite complex and especially come at the cost of a very large increase in GPU global memory access, which massively slows down the calculation. However, there is also a second motivation to splitting Feynman diagrams, namely making the code easier to build and manage for very complex physics processes with very large numbers of diagrams. As described in this section, my work in this direction did achieve some interesting progress, such as the possibility to compute $2\!\rightarrow\!6$ and $2\!\rightarrow\!7$ processes which were previously non-computable in the CUDACPP software.
In the following subsections, I describe my second batch of three sets of kernel splitting changes, which I now also consider ready to be integrated in new production releases of CUDACPP: (ihel4) one kernel per Feynman diagram in each helicity stream; (ihel4p1) ditto, with optional CUDA graph orchestration; (ihel5) one kernel per diagram group in each helicity stream, with separate source code files; (ihel6) a single kernel in each helicity stream, processing all diagram groups as device functions. The results of the intermediate ihel4 step and those of the final ihel6p2 step on $gg\!\rightarrow\!t\bar{t}$ production with 0, 1, 2 or 3 extra final state gluons are given in Figures 4 and 5 for NVidia and AMD GPUs: these plots show, in particular, that the achieved throughputs in the ihel6p2 software are completely equivalent to those of the earlier ihel3p1 software, when the code is generated without diagram splitting, i.e. with a single kernel processing all Feynman diagrams as in ihel3p1. The main interest of ihel6p2, conversely, comes from the more complex $gg\!\rightarrow\!t\bar{t}gggg$ and $gg\!\rightarrow\!t\bar{t}ggggg$ with 4 and 5 extra final state gluons, the results for which are presented at the end of this section.
Figure 10: Throughputs (ME/s) as a function of grid size for an NVidia V100 GPU (node itscrd90 at CERN). Higher is better. The 2 plots correspond to throughputs derived by executing the check.exe application through 1 (left) or 10 (right) cycles of a GPU grid. The number of threads per block is fixed to 32 (warp size); the grid size is varied by changing the number of blocks. Each plot compares different scenarios considered in this paper: (ihel3) all Feynman diagrams in a single kernel; (ihel4) Feynman diagrams as individual kernels, without CUDA graphs; (ihel4p1) Feynman diagrams as individual kernels, with CUDA graphs.
4.1 Feynman diagrams as individual kernels (“ihel4”)